OpenStudy (anonymous):

Find the exact area of the circle with the given diameter. d = 12.6 cm A. 39.69 cm2 B. 158.76 cm2 C. 158.76π cm2 D. 39.69π cm2

OpenStudy (nincompoop):

you know the formula of a circle, yes?

OpenStudy (anonymous):

yes

OpenStudy (richyw):

\[a_\text{circle}=\pi r^2\]\[\qquad=\pi\left(\frac{d}{2}\right)^2\]\[\qquad=\pi\frac{d^2}{4}\]

OpenStudy (anonymous):

is it B

OpenStudy (richyw):

not quite, you forgot to multiply it by pi

OpenStudy (nincompoop):

if you know the formula for the area of a circle, then why are you asking us if the answer is B? how about you show us your solution eh